# Self-Supervised Learning and Opportunistic Inference for Continuous Monitoring of Freezing of Gait in Parkinson’s Disease

### v1

@@
Abstract
Parkinson’s disease (PD) is a progressive neurological disorder that impacts the quality of life significantly, making in-home monitoring of motor symptoms such as Freezing of Gait (FoG) critical. However, existing symptom monitoring technologies are power-hungry, rely on extensive amounts of labeled data, and operate in controlled settings. These shortcomings limit real-world deployment of the technology. This work presents LIFT-PD [^1] The source code is available at: https://github.com/shovito66/LIFT-PD, a computationally-efficient self-supervised learning framework for real-time FoG detection. Our method combines self-supervised pre-training on unlabeled data with a novel differential hopping windowing technique to learn from limited labeled instances. An opportunistic model activation module further minimizes the power consumption by selectively activating the deep learning module only during active periods. Extensive experimental results show that LIFT-PD achieves a 7.25% increase in precision and 4.4% improvement in accuracy compared to supervised models while using as low as 40% of the labeled training data used for supervised learning. Additionally, the model activation module reduces inference time by up to 67% compared to continuous inference. LIFT-PD paves the way for practical, energy-efficient, and unobtrusive in-home monitoring of PD patients with minimal labeling requirements.
{IEEEkeywords}
@@ Abstract
1 Introduction
[…]
@@ 1 Introduction
Parkinson’s disease (PD) is a progressive neurological disorder affecting 7–10 million people worldwide, significantly impacting their quality of life. Freezing of Gait (FoG), a transient inability to produce effective steps during walking [1], is one of the most debilitating symptoms of PD, leading to poor mobility, increased risk of falls and injuries, and reduced quality of life. While treatments like levodopa can sometimes reduce the severity of freezing episodes, their effectiveness is often incomplete and variable, diminishing over time [2]. Compensatory treatments such as on-demand cueing require patient or companion initiation, which can be challenging in time-sensitive or anxiety-provoking situations that trigger freezing [3]. Identifying FoG events and, more importantly, the time leading up to a FOG event could result in early deployment of on-demand cues to help reduce the severity of a freezing event [4].
@@
In recent years, wearable technology and machine learning (ML) algorithms have emerged as promising tools for continuous monitoring and management of PD symptoms, including gait disturbances and tremors [5, 6, 7] However, real-world deployment of these technologies for continuous monitoring and intervention has remained a significant research challenge. One major hurdle is the scarcity of labeled data essential for training robust ML models, particularly in PD, where symptom manifestation is highly individualized. Data annotation requires considerable time and expertise, limiting the availability of accurate and diverse datasets [8].
@@
Detecting FoG using a patient-independent model is a complex task due to significant inter- and intra-variability in patients’ gait patterns. Previous research on FoG detection relied on multiple sensors, extensive feature engineering, patient-specific data collection, and model retraining [9, 10, 11]. To address these challenges, we present an innovative label-efficient, patient-independent, and robust self-supervised learning framework, LIFT-PD (Label-efficient In-home Freezing-of-gait Tracking in Parkinson’s Disease), for detecting FoG events in real-time. Our extensive results demonstrate that the proposed self-supervised model achieves comparable performance to baseline supervised models, despite using much fewer labeled instances and in an unsupervised fashion. Furthermore, we devise a novel opportunistic-based lightweight algorithm that reduces the complexity of model execution and processing time, paving the way for implementation in a stand-alone wearable system for real-time symptom monitoring.
Table 1: Summary of Recent FoG Detection Works. FI:Freezing Index, SP: Stride Peak, STD:Standard Deviation, SD: Subject Dependent, SI: Subject Independent, LOOCV: Leave-One-Out Cross-Validation
Study Method Sensor (Number) (Position) Preprocessing/ Extracted Features Validation Contribution
 [9] CNN and Transformer (Supervised) Accelerometer (1) (Left Waist) Resample, Filtering, FFT, FI LOSO CV (SI) Transformer and CNN based methodologies for FoG detection, along with a clustering approach for FoG episode analysis
 [12] Multihead CNN (Supervised) $\times$ Hold Out (SI) Light algorithm for real-time FoG detection using raw sensors data.
 [5] Supervised ML IMU (2) (Shin) Time & Frequency Domain Features 10 Fold CV, LOSO train-test 70-30% Detect and predict FoG considering the impact of dopaminergic therapy on performance.
 [11] Semi Supervised 3 Layers IMU (1) (Ankle) Filtering, 3 Features (FI, SP, STD) LOSO (SD) Adapting FoG classifier’s parameters in real-time with unlabeled data
 [10] Supervised, transfer learning IMU (2) (Left-Right Ankle) Filtering, Data Augmentation LOOCV (SI) One-class classifier for FoG detection using only normal gait data
2 Related Work
Prior work can be summarized into two categories including (1) self-supervised learning (SSL) for motion analysis; and (2) automatic detection of FoG events.
The SimCLR SSL technique was originally designed for computer vision tasks,  [13], but was later adapted for human activity recognition (HAR) using a transformer-based encoder method [14]. Although MLPs are not popular for vision or language datasets, they are effective in making predictions on tabular health data  [15, 16]. Previous work also introduced transfer learning [16], data augmentation technique and resampling, which are efficient in scenarios with limited labeled data [17]. The efficacy of self-supervised learning in medical time series analysis has been demonstrated, emphasizing its role in data augmentation and contrastive pair formation [18].
The detection of FoG events in PD has been the subject of extensive research, often involving multimodal datasets. In particular, prior work explored the effectiveness of different sensor modalities such as gait acceleration, electroencephalogram (EEG), electromyography (EMG), and skin conductance (SC) in FoG detections [19]. Furthermore, researchers used an LSTM-based model combining acceleration and synthetic EEG data [20].
Table 1 summarizes recent PD studies on FoG detection. LIFT-PD overcomes key limitations of these studies, such as limited generalizability, reliance on extensive labeled data, and high computational complexity. In contrast, our proposed method, LIFT-PD, addresses these limitations by introducing a self-supervised learning framework tailored for real-time FoG detection using a single accelerometer sensor.
3 LIFT-PD System Design
@@
Figure 1: Self-Supervised training pipeline for real-time FoG detection. Pretrain: Encoder reconstructs masked signal segments, outputting predicted values $h_{i}$ for missing data. Fine-tuning: Encoder weights are frozen while the MLP is optimized using cross-entropy loss on labeled data. End-to-End Pipeline: The trained model is integrated into a real-time system, where model activation module (MAM) selectively activates the computationally intensive FoG detection model for energy-efficient, long-term monitoring.
@@
Training robust deep learning (DL) models typically requires large amounts of labeled data, which can be challenging to obtain, especially for tasks like freezing of gait (FoG) detection in elderly. To address this issue, LIFT-PD employs self-supervised learning to utilize unlabeled data during training. Another challenge that we face while designing ML models for FoG detection is that FoG events are sparse, resulting in an imbalanced dataset with a significantly lower proportion of FoG episodes compared to non-FoG activities. To mitigate this challenge, we incorporate the Differential Hopping Windowing (DHW) Technique during data preprocessing, which applies variable overlaps for FoG and non-FoG instances, thereby enhancing the model’s ability to learn from the underrepresented class. Finally, an opportunistic-based lightweight algorithm is introduced to reduce execution complexity, allowing for implementation in stand-alone wearable devices (Section 3.2).
3.1 Problem Modeling by Self-Supervised Learning (SSL)
@@
LIFT-PD uses SSL for FoG detection in two steps: learning contextual representations from raw signals using a 1D CNN model, followed by performing the downstream FoG detection task. In this paper, raw accelerometer signals are used as physiological signals and the downstream task is binary ‘Freezing of Gait’ detection. In the following subsections, we describe in detail the two main components of our SSL approach:
3.1.1 Pretext Task (Signal Reconstruction)
@@
During pre-training, a 3-layer neural network model with an encoder is trained using SSL. Random segments of fixed length $m$ are masked from each window, $x_{w_{i}}$ by replacing them with zeros. The model is then trained to predict these masked values. as shown in Fig. 1 (Part 1). This whole training process does not need any labeled data. The masked signal data $\hat{X}_{W}\in\mathbb{R}^{N\times T^{\prime}}$, is defined by $\hat{X}_{W}=mask(X_{W},m)$ where $\hat{x}_{w_{i}}\in\hat{X}_{W}$
@@
The encoder $f_{\theta}:\mathbb{R}^{N\times T^{\prime}}\rightarrow\mathbb{R}^{N\times D}$ reconstructs the original signal from the masked windows, producing a lower dimensional hidden representation $H\in\mathbb{R}^{N\times D}$, where $h_{i}\in\mathbb{R}^{D}\leftarrow f_{\theta}(\hat{x}_{w_{i}})$. These are passed through a fully connected network $g_{\psi}(\cdot)$ to predict the masked segments represented as $\hat{x}=g_{\psi}(h_{i})$. The encoder minimizes the reconstruction loss $L_{\theta}$ using mean squared error (MSE) defined as $L_{\theta}=\frac{1}{N_{m}}\sum_{j=1}^{N_{m}}(\hat{x}_{j}-x_{p(j)})^{2}$, where $N_{m}$ is the total number of masked points, $\hat{x}_{j}$ is the prediction for the $j-$th masked point, and $x_{p(j)}$ is the original input value, whose position is $p(j)$ in the segmented signal windows.
3.1.2 Downstream Task (FoG Classification)
@@
After pre-training, a Multilayer Perceptron (MLP) $g_{\phi}:\mathbb{R}^{N\times D}\rightarrow Y$, is introduced on top of the encoder for binary FoG detection. The MLP maps the learned embeddings $\mathbf{H}$ to FoG or non-FoG labels $\mathbf{Y}$, represented as $\hat{Y}=g_{\phi}(H)=g_{\phi}(f_{\theta}(X_{W}))$, where $f_{\theta}$ is the pre-trained encoder with parameters $\theta$.
@@
The fine-tuning step is supervised, using a small amount of labeled data, with the encoder weights frozen and only the MLP layers trained with a lower learning rate as shown in Fig. 1 (Part 2). We minimize the binary cross-entropy loss $L_{ce}{(\theta,\phi)}$ for the downstream task defined as $L_{ce}(\theta,\phi)=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\sigma(\hat{y}_{i}))+(1-y_{i})\log(1-\sigma(\hat{y}_{i}))\right]$, where N is the total number of windows, $y$ is an indicator variable (1 for FoG, 0 for non-FoG), $\sigma$ is the sigmoid function, and $\hat{y}\in\hat{Y}$ is the output of the classifier $g_{\phi}(h_{i})$.
3.2 Opportunistic Inference

### v2

@@
Abstract.
Parkinson’s disease (PD) significantly affects patients’ quality of life through debilitating motor symptoms, such as Freezing of Gait (FoG). Continuous, in-home monitoring of FoG is essential for timely clinical intervention but remains challenging due to high power consumption, annotation cost, and the controlled environments required by current wearables. We introduce LIFT-PD [^1] The source code is available at: https://github.com/shovito66/LIFT-PD, a novel self-supervised learning (SSL) framework for real-time, patient-independent FoG detection that uniquely utilizes a single waist-worn accelerometer—an approach traditionally considered less optimal due to weaker gait signatures. LIFT-PD leverages SSL on unlabeled data collected from uncontrolled, real-world settings and employs a novel Differential Hopping Windowing Technique (DHWT) to address gait variability and dataset imbalance. Additionally, an opportunistic inference module selectively activates the deep learning model only during patient movement, significantly reducing power consumption and enabling continuous monitoring (>>48 hours). Experimental results show that LIFT-PD achieves a 7.25% increase in precision and 4.4% improvement in accuracy compared to supervised and semi-supervised baseline models while requiring approximately 40% fewer labeled training samples. Evaluations across diverse patient characteristics—including severity, medication state, age, and gender—confirm the model’s robustness and clinical applicability, positioning LIFT-PD as a practical, energy-efficient, and scalable solution for continuous real-world FoG monitoring in PD.
@@ Abstract.
This work was supported in part by the National Science Foundation under grant CNS-2227002 as well as the ASU College of Health Solutions. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding organizations.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: HEALTHjournalvolume: TBDjournalnumber: TBDarticle: TBDisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Self-supervised learningccs: Computing methodologies Semi-supervised learning, Neural networks, Wearable Sensors
1. Introduction
@@
Parkinson‘s disease (PD) is a progressive neurological disorder affecting 7–10 million people worldwide, significantly impacting their quality of life (soumma2024mds). Freezing of Gait (FoG), a transient inability to produce effective steps during walking (nieuwboer2013characterizing), is one of the most debilitating symptoms of PD, leading to poor mobility, increased risk of falls and injuries, and reduced quality of life. While treatments like levodopa can sometimes reduce the severity of freezing episodes, their effectiveness is often incomplete and variable, diminishing over time (zhang2016freezing). Compensatory treatments such as on-demand cueing require patient or companion initiation, which can be challenging in time-sensitive or anxiety-provoking situations that trigger freezing (cosentino2023one). Identifying FoG events and, more importantly, the time leading up to a FoG event could result in early deployment of on-demand cues to help reduce the severity of a freezing event (jadhwani2023review).
@@
In recent years, wearable technology and machine learning (ML) algorithms have emerged as promising tools for continuous monitoring and management of PD symptoms, including gait disturbances and tremors (borzi2021prediction; bikias2021deepfog; SilvadeLima2017FreezingOG) However, real-world deployment of these technologies for continuous monitoring and intervention has remained a significant research challenge. One major hurdle is the scarcity of labeled data essential for training robust ML models, particularly in PD, where symptom manifestation is highly individualized. Data annotation requires considerable time and expertise, limiting the availability of accurate and diverse datasets (peppes2023foggan).
@@
Detecting FoG using a patient-independent model is a complex task due to significant inter- and intra-variability in patients’ gait patterns. Previous research on FoG detection relied on multiple sensors, extensive feature engineering, patient-specific data collection, and model retraining, limiting large scale adoption of wearable technologies for long-term health monitoring (sigcha2022improvement; naghavi2021towards; mikos2017real). To address these challenges, we present an innovative label-efficient, patient-independent, and robust self-supervised learning framework, LIFT-PD (Label-efficient In-home Freezing-of-gait Tracking in Parkinson’s Disease), for detecting FoG events in real-time.
The main contributions of this work are:
  1. (1)
    Label-Efficient Learning: LIFT-PD uses a self-supervised learning approach, allowing the model to learn from unlabeled data with minimal labeled instances required for fine-tuning. This significantly reduces the need for extensive data annotation, making it more practical for real-world deployment where labeled data is often limited.
  2. (2)
    Handling Variability and Imbalanced Data: Our Differential Hopping Windowing Technique (DHWT) effectively handles inter and intra-patient gait variations and data imbalance issues.
  3. (3)
    Robust Performance: Despite using significantly fewer labeled data points, LIFT-PD achieves comparable performance to fully supervised approaches.
  4. (4)
    Energy Efficiency: LIFT-PD includes an opportunistic inference module that activates the model only during active movement periods, significantly reducing computational load and power consumption, paving the way for implementation in a stand-alone wearable system for real-time symptom monitoring.
  5. (5)
    Practical Deployment: By using a single triaxial accelerometer and processing data in real-time without extensive preprocessing, LIFT-PD offers a practical, scalable solution for continuous, in-home real-time monitoring of PD patients, facilitating seamless integration into everyday healthcare systems.
These contributions make LIFT-PD a scalable, energy-efficient, and patient-independent solution for real-time FoG detection, with significant potential for improving Parkinson’s disease management through continuous and accessible monitoring.
@@
Figure 1. Self-Supervised training pipeline for real-time FoG detection. Pretrain: Encoder reconstructs masked signal segments, outputting predicted values $h_{i}$ for missing data. Fine-tuning: Encoder weights are frozen while the MLP is optimized using cross-entropy loss on labeled data. End-to-End Pipeline: The trained model is integrated into a real-time system, where model activation module (MAM) selectively activates the computationally intensive FoG detection model for energy-efficient, long-term monitoring.
Table 1. Summary of Recent FoG Detection Works. FI:Freezing Index, SP: Stride Peak, STD:Standard Deviation, SD: Subject Dependent, SI: Subject Independent, LOOCV: Leave-One-Out Cross-Validation
Study
Method
Sensor (Number)
Preprocessing/
Extracted Features
Validation
Contribution
 (sigcha2022improvement) CNN, Transformer (Supervised)
Accelerometer (1)
(Left Waist)
Resample, Filtering, FFT, FI LOSO CV (SI) Transformer and CNN based methodologies for FoG detection, along with a clustering approach for FoG episode analysis
 (10.1016/j.artmed.2022.102459) Multihead CNN (Supervised) $\times$ Hold Out (SI) Light algorithm for real-time FoG detection using raw sensors data.
 (borzi2021prediction) Supervised ML IMU (2) (Shin) Time & Frequency Domain Features 10 Fold CV, LOSO train-test 70-30% Detect and predict FoG considering the impact of dopaminergic therapy on performance.
 (mikos2017real) Semi Supervised 3 Layers
IMU (1)
Filtering, 3 Features
(FI, SP, STD)
LOSO (SD) Adapting FoG classifier’s parameters in real-time with unlabeled data
Supervised,
transfer learning
IMU (2)
(Left-Right Ankle)
Filtering,
Data Augmentation
LOOCV (SI)
One-class classifier for FoG detection using only normal gait data
2. Related Work
The detection of FoG events in PD has been the subject of extensive research, often involving multimodal datasets. In particular, prior work explored the effectiveness of different sensor modalities such as gait acceleration, electroencephalogram (EEG), electromyography (EMG), and skin conductance (SC) in FoG detections (zhang2022multimodal). Furthermore, researchers used an LSTM-based model combining acceleration and synthetic EEG data (guo2022high). Previous work also introduced transfer learning (10331481), data augmentation technique and resampling, which are efficient in scenarios with limited labeled data (wang2022sensor). We can broadly categorize these prior works based on label usage into supervised learning approaches for motion analysis and weakly labeled approaches, including semi-supervised and self-supervised learning, for FoG detection.
Supervised methods for FoG detection often rely on large amounts of labeled data, which is a significant limitation in real-world applications. These methods typically use machine learning (ML) algorithms such as CNNs, transformers, and LSTMs (guo2022high) combined with sensor modalities like accelerometers and IMUs. Multi-head CNN (10.1016/j.artmed.2022.102459) and CNN-Transformer models (sigcha2022improvement) have shown good performance in detecting FoG in PD patients, but they face challenges in generalization and high computational requirements, which limit their feasibility for real-time wearable monitoring. While these models provide high detection rates, their reliance on extensive feature engineering (borzi2021prediction; koltermann2024gait), a large number of training samples, and significant computational resources makes them impractical for wearable, real-time monitoring. Additionally, the models struggle with high false-positive rates, which can reduce their effectiveness in real-world settings where minimizing such errors is crucial for patient safety and usability. Transfer learning techniques, such as those used in the One-Class Classifier (naghavi2021towards), help improve detection in normal conditions but fail to adapt to the dynamic and complex nature of FoG episodes, limiting their effectiveness in real-world applications. Table 1 summarizes recent PD studies on FoG detection.
To address the challenge of insufficient labeled data, some studies have turned to semi-supervised learning (semi-SL), using a combination of labeled and unlabeled data for model training. Mikos et al. (mikos2017real) uses a three-layer network with an IMU at the ankle, incorporating filtering and feature extraction. While the approach shows improvement over purely supervised models, it still requires a considerable amount of labeled data and suffers from issues related to the inability to generalize effectively across different PD patient profile.
The application of self-supervised learning (SSL) to time-series data, particularly for human activity recognition (HAR), is a newer area of research. Unlike supervised and semi-supervised methods, SSL frameworks like SimCLR (chen2020simple) focus on learning representations without relying on labeled data, making them ideal for scenarios with limited annotations. Originally, it was designed for computer vision tasks, but was later adapted for HAR using a transformer-based encoder method (khaertdinov2021contrastive). However, the use of SSL for FoG detection remains limited due to challenges in effectively balancing the dataset and ensuring model generalization to diverse populations. The efficacy of self-supervised learning in medical time series analysis has been demonstrated, emphasizing its role in data augmentation and contrastive pair formation (liu2023self).
While numerous studies have proposed effective methods for FoG detection, most rely on either extensive labeled data or computationally expensive models that are unsuitable for real-time, wearable deployment. These methods often depend on data collected in controlled environments, limiting their practical utility for everyday monitoring. Additionally, existing approaches struggle with imbalanced datasets, where FoG events are underrepresented and have limited generalization across diverse patient profiles, such as varying severities, ages, and genders. In contrast, LIFT-PD introduces a self-supervised framework that addresses these challenges by leveraging a Differential Hopping Windowing Technique (DHWT) for handling class imbalances during training. By using a minimal amount of labeled data and combining Opportunistic Inference Modules to optimize power consumption, LIFT-PD achieves real-time FoG detection with significantly reduced computational overhead. LIFT-PD’s use of a single triaxial accelerometer and real-time data processing without extensive preprocessing makes it highly practical for wearable devices, unlike traditional models that rely on multiple sensors and are too resource-intensive for continuous monitoring.
3. LIFT-PD System Design
@@
Training robust deep learning (DL) models typically requires large amounts of labeled data, which can be challenging to obtain, especially for tasks like freezing of gait (FoG) detection in elderly. To address this issue, LIFT-PD employs self-supervised learning to utilize unlabeled data during training. Another challenge that we face while designing ML models for FoG detection is that FoG events are sparse, resulting in an imbalanced dataset with a significantly lower proportion of FoG episodes compared to non-FoG activities. To mitigate this challenge, we incorporate the ’Differential Hopping Windowing Technique (DHWT) during data preprocessing, which applies variable overlaps for FoG and non-FoG instances, thereby enhancing the model’s ability to learn from the underrepresented class. Finally, an opportunistic-based lightweight algorithm is introduced to reduce execution complexity, allowing for implementation in stand-alone wearable devices (Section 3.2).
3.1. Problem Modeling by Self-Supervised Learning (SSL)
@@
LIFT-PD uses SSL for FoG detection in two steps: learning contextual representations from raw signals using a 1D CNN model, followed by performing the downstream FoG detection task. In this paper, raw accelerometer signals are used as physiological signals and the downstream task is binary ‘Freezing of Gait’ detection.
The first stage-SSL pre-training—is performed through a pretext task of masked-signal reconstruction, where the model learns meaningful temporal representations from unlabeled data. In the following subsections, we describe in detail the two main components of our SSL approach:
3.1.1. Pretext Task (Signal Reconstruction)
@@
During pre-training, a 3-layer neural network model with an encoder is trained using SSL. Random segments totaling 20% of the window length are masked from each window ($x_{w_{i}}$) by replacing them with zeros. The model is then trained to predict these masked values as shown in Fig. 1 (Part 1). Preliminary experiments comparing masking ratios of 10%, 20%, and 30% showed that 20% achieved the best trade-off—posing enough reconstruction difficulty to learn robust representations while preserving sufficient temporal context for meaningful feature learning. This whole training process does not need any labeled data. The masked signal data $\hat{X}_{W}\in\mathbb{R}^{N\times T^{\prime}}$, is defined by $\hat{X}_{W}=mask(X_{W},m)$ where $\hat{x}_{w_{i}}\in\hat{X}_{W}$ and $m$ represents the masking ratio (0.2 in our implementation).
@@
The encoder $f_{\theta}:\mathbb{R}^{N\times T^{\prime}}\rightarrow\mathbb{R}^{N\times D}$ reconstructs the original signal from the masked windows, producing a lower dimensional hidden representation $H\in\mathbb{R}^{N\times D}$, where $h_{i}\in\mathbb{R}^{D}\leftarrow f_{\theta}(\hat{x}_{w_{i}})$. These are passed through a fully connected network $g_{\psi}(\cdot)$ to predict the masked segments represented as $\hat{x}=g_{\psi}(h_{i})$. The encoder minimizes the reconstruction loss $L_{\theta}$ using mean squared error (MSE) defined as:,
(1) $L_{\theta}=\frac{1}{N_{m}}\sum_{j=1}^{N_{m}}$ $(\hat{x}_{j}-x_{p(j)})^{2}$
where $N_{m}$ is the total number of masked points, $\hat{x}_{j}$ is the prediction for the $j$-th masked point, and $x_{p(j)}$ is the original input value, whose position is $p(j)$ in the segmented signal windows. The objective is to find the optimal parameters $\theta^{*}$ that minimize this $L_{\theta}$ loss, which we can formally define as below:
$\theta^{*}=\arg\min_{\theta}L_{\theta}$
3.1.2. Downstream Task (FoG Classification)
@@
After pre-training, a Multilayer Perceptron (MLP) $g_{\phi}:\mathbb{R}^{N\times D}\rightarrow Y$, is introduced on top of the encoder for binary FoG detection. The MLP maps the learned embeddings $\mathbf{H}$ to FoG or non-FoG labels $\mathbf{Y}$, represented as $\hat{Y}=g_{\phi}(H)=g_{\phi}(f_{\theta}(X_{W}))$, where $f_{\theta}$ is the pre-trained encoder with parameters $\theta$ and $X_{W}$ is the segmented signal data.
@@
This fine-tuning step is supervised, using a small amount of labeled data. Initially, the encoder $f_{\theta}$ is initialized with the pre-trained weights $(\theta^{*})$, while MLP $g_{\phi}$ is randomly initialized. During the fine-tuning the encoder weights remains frozen and only the MLP layers are trained with a low learning rate as shown in Fig. 1 (Part 2). For the downstream FoG detection task, we minimize the binary cross-entropy loss $L_{ce}{(\theta,\phi)}$:
(2) $L_{ce}{(\theta,\phi)}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\sigma(\hat{y}_{i}))+(1-y_{i})\log(1-\sigma(\hat{y}_{i}))\right]$
where N is the total number of windows, $y$ is an indicator variable (1 for FoG, 0 for non-FoG), $\sigma$ is the sigmoid function, and $\hat{y}\in\hat{Y}$ is the output of the classifier $g_{\phi}(h_{i})$ for the $i$’th window of embedding $h_{i}$.
3.2. Opportunistic Inference